OpenStudy (anonymous):

Write the solution as a set of ordered pairs given the equation y = 7.55x + 300, when x = 0, 100, 160.

jimthompson5910 (jim_thompson5910):

When x = 0, y = 7.55(0)+300 = 300 So when x = 0, y = 300 So the first ordered pair is (0,300) ------- When x = 100, y = 7.55(100)+300 = 755+300 = 1055 So when x = 100, y = 1055 So the second ordered pair is (100,1055) ------- When x = 160, y = 7.55(160)+300 = 1208+300 = 1508 So when x = 160, y = 1508 So the third ordered pair is (160,1508)